# Auto-Adaptive PINNs with Applications to Phase Transitions

### v3

@@ 2.2. Energy-Adaptive Sampling
For our simulations of the Allen-Cahn equation, we will use the heuristic provided in section 1.2.3 as our sampling distribution. That is, we will sample in proportion to the pointwise energy density of the function approximation. We additionally incorporate a weight in front of the adaptive term of the loss. This can be interpreted as weighting the relative importance for the separation of scales in time, while the sampling can be viewed as weighting the relative importance or difficulty of capturing different regions in space. In fact, this weight is exactly $\sqrt{\gamma_{2}/\gamma_{1}}$, which dictates the separation of timescales. This weight can be seen as setting the relative importance of the uniform distribution as compared to the adaptive distribution.
@@
Notably, we do not allow the spatially adaptive points to move in time. Instead we fix the sample points in time, and use the Metropolis-Hastings algorithm to move the points only in space to represent the energy density at that fixed time. This is because if allowed to move in time, points will coalesce at the beginning of the time interval (as the total energy of the system decreases according to equation 1.5). Exploring time dependent adaptive densities in an obvious direction to be explored in future work.
3. Implementation
@@ 3.1. Time Slicing
This method was introduced in [WZ] and involves the gradual expansion of the time domain in discrete steps. The idea is to reduce the time complexity of the problem by first greatly restricting the time domain until it is palatable to the network. Once the problem is learned on that smaller interval, the interval can be increased. This allows for the network to only require learning a small time interval at any given time time, which is generally easier for training. This method was tested very thoroughly with a large degree of success. However, for higher-dimensional and less regular systems further issues arise.
One of the these issues is commonly refereed to as ‘Catastrophic Unlearning’ in the wider Machine Learning community, though use of this term is not commonly used in the discussion of PINNs. This issue is common in traditional data driven contexts where networks are trained to do one task and then trained to do a second separate task. In this instance the network may unlearn the original task in favor of learning the second task. In the context of PINNs this issue manifests when a network which is well trained on some sub-interval of the entire time domain, is trained on a separate or expanded time interval. As a particular example, a PINN trained on the time interval $[0,.1]$ may drastically lose accuracy when trained on $[0,.2]$ since simulating from .1 to .2 can be considered a ‘new task.’ Not only will the simulation not perform accurately on the new time window of .1 to .2, but it will also lose the accuracy it had on the original window of 0 to .1. Of course simulating accurately on $[0,.1]$ is essential to simulating on $[.1,.2]$, so in the setting of PINNs the problem is even further accentuated.
The authors of [WZ] were aware of this issue, so they presented an alternative method that they call ‘Time Slicing II.’ This method involves the training of a separate network on each time slice rather than using the gradual expansion of training time on a single network. This does resolve the issue of catastrophic unlearning, since no network is trained to perform multiple tasks. However, it is somewhat undesirable as it requires many networks, a true time discretization, and can also dramatically increase cost in training time and number of network parameters. For these reasons we manually edit the learning rate as we evolve through time slices, which is discussed next.
@@ 3.2. Learning Rate Schedule
We additionally use a learning rate scheduler, which updates the learning rate as the training process progresses. We use a fairly simple scheduler which updates the learning rate only based off the current time slice being trained. This helps to combat the catastrophic unlearning described in the section above. We decrease the learning rate roughly linearly as the progress through the time-slices. More sophisticated methods of choosing the learning rates are constantly being studied [lrannealing, softadapt], however for simplicity and ease of implementation we choose to use only to use this simple scheme.
@@ 3.3. Residual Adaptive Sampling
A commonly employed method to increase training efficacy is Residual Adaptive sampling, first proposed in [orig_resadapt]. This technique involves sampling the domain in proportion to the current pointwise value of the residual. This leads to decreasing the loss very efficiently, as has been thoroughly explored in [resadapt], and seen in many other papers including [WZ]. In our numerical experiments, we will use this method as a test to compare our own methods against.
@@ 3.4. Minibatching
3.5. Details of Metropolis Hastings
In this section we describe in more detail the precise nature of the Metropolis-Hastings Procedure we use in our experiments. In general, we err on the side of using too many iterations than too few. When initializing the adaptive points on each time slice, we perform $10,000$ iterations. This is considerably more than necessary, but nearly guarantees the convergence to an appropriate distribution. We also step the adaptive points at the conclusion of each training epoch (not after each minibatch). In this step, we perform $200$ iterations.
As a proposal distribution, we use a normal distribution centered around each point. The standard deviation for this distribution is chosen as $\sqrt{\gamma_{1}}$, which is the separation of spatial scales. The standard deviation is then updated each iteration to target an acceptance rate between .2.2 and .6.6.
@@ 3.6. Latin Hypercube Sampling
For uniform sampling in our domain, we use Latin Hypercube sampling. This method reduces variance in sampling by partitioning the domain into a grid of $N^{d}$ hypercubes, where $N$ is the desired number of sample points and $d$ is the number of spatial dimensions. Then hypercubes are sampled randomly so that the selected hypercubes are orthogonal. This yields a selection of $N$ hypercubes in the grid, with no selected hypercube in the same row or column as any other. Once this orthogonal set of hypercubes is selected, one point within each is chosen randomly from a uniform distribution on that hypercube. These individual sample points combine to form a representative sample of $N$ points from the domain. Further explanation can be found [hypercube]. This method guarantees that the uniform sample will be representative of the entire domain. This sampling also greatly reduces variance, and thus is less reliant on the convergence yielded by the law of large numbers. This especially helps to reduce inconsistencies of sampling too few points, and makes the training more robust.
@@ 4. Numerical Results
The following numerical benchmarks are taken from [WZ]. In the examples they provide, there is a clear concentration of the error in the regions of high energy. As such, this is an ideal setting for the testing of the improvements given by our new method.
For each example, we compare three methods. First, we run a high fidelity finite difference to yield an ‘exact’ solution. The second method is Residual Adaptive sampling. Here we use the Metropolis Hastings Algorithm to sample in proportion to the loss and combine these with points sampled uniformly as in [resadapt]. Additionally this implementation uses all of the techniques described in the implementation section, including time slices, learning rate scheduling, minibatching, latin hypercube sampling, and initial condition weights. This is meant to approximate current state of the art techniques. Finally we implement our energy adaptive method described in section 3.
@@
(4.1) $\gamma_{1}=1e-4,\hskip 5.69046pt\gamma_{2}=5,\hskip 5.69046ptu_{0}(x)=x^{2}\cos(\pi x)$
@@
Our network uses 6 fully connected layers of 128 nodes. We use the hyperbolic tangent activation function for all interior layers. For training we use $10,000$ total collocation points in the domain. We train the network until final time $t=1$, with time slices taken in increments of .1.1. On each time slice, we run 100 epochs of ADAM training with minibatches of size $40$ split proportionally according to the adaptive point ratio, $\lambda$. As we progress through time slices, we additionally change the learning rate. We use learning rate $10^{-3}$ until final time .3, at which point we switch to $5*10^{-4}$. At final time .5.5, we decrease to learning rate $10^{-4}$. At .7.7 we decrease to $5*10^{-5}$, and finally at .9.9 we decrease to $10^{-5}$.
We perform a parameter sweep on $\lambda$, the proportion of adaptively chosen points vs uniformly chosen points. The results for this are found in Figure 1. We can see from this plot that the energy adaptive method performs best with a $\lambda$ value of about .6.6, while residual adaptive functions about equivalent for $\lambda$ between .6.6 and .9.9. For further discussion we consider $\lambda=.6$ so as to more directly compare with the energy adaptive method. Notice also the consistent improvement of the energy adaptive method in both error measures.
@@
To verify that both methods are decreasing loss effectively, we display the loss across each epoch in Figure 2. It is verified here that each method is successfully evolving according to their respective losses. This plot also informs our decision to gradually anneal the learning rate.
@@
Figure 3. Time slices from networks which have completed the described training. We present time slices from the the residual adaptive method and the energy adaptive method at times 0, .25.25, .5.5, and $1$.
@@
Relative $L^{2}$ at $T=1$ 4.09e-02 1.50e-02
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error 2.31e-01 7.96e-02
@@
(4.2) $\gamma_{1}=1e-4,\hskip 5.69046pt\gamma_{2}=4,\hskip 5.69046ptu_{0}(x)=x^{2}\sin(2\pi x)$
@@
Relative $L^{2}$ at $T=1$ 2.33e-02 6.87e-03
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error 1.15e-01 3.20e-02
@@
Finally we experiment with the 2D Example given in [WZ]. This is characterized by the parameters
(4.3) $\gamma_{1}=\lambda\epsilon^{2},\hskip 8.5359pt\gamma_{1}=\lambda,\hskip 8.5359ptu_{0}(x)=\tanh\left(\frac{.35-\sqrt{(x-.5)^{2}+(y-.5)^{2}}}{2\epsilon}\right)$
@@
Relative $L^{2}$ at $T=10$ $T=9$ 1.52e-01 3.56e-02
Relative $L^{2}$ at $T=10$ $T=10$ 3.53e-01 1.93e-01
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error $T=9$ 1.73e+$0$ 9.79e-01
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error $T=10$ 1.98e+$0$ 1.90e+$0$
@@
In general there are several methods which can be used to reduce the problems posed by catastrophic unlearning. The two most prominent of which would be the employment of an xPINN [xPINN] or the use of exponential time sampling [ExpTime]. These directions are known to be effective, but require a post-hoc knowledge of where the unlearning occurs. For our analysis here, we are trying to confine our approach only to methods which can be applied without prior knowledge of the location of the unlearning. We also believe that an intelligently chosen heuristic employed via auto-adaptive PINNs could efficiently resolve this issue. However, we leave the exploration of this as a future direction.
@@ 5. Conclusions and Future Directions
Analytic verification in the form of a proof is also desirable for the energy adaptive method, perhaps in a similar manner to the proofs in [ExpTime]. If well understood, this could also motivate choices of sampling densities for other difficult problems.

### v4

@@ 2.1. Energy-Adaptive Sampling
For our simulations of the Allen-Cahn equation, we will use the network dependent heuristic provided in section 1.2.3 as our sampling density. That is, we will sample in proportion to the pointwise energy density of the function approximation.
Explicitly, our target density for the Metropolis Hastings Algorithm will be given (up to scaling) by:
(16) $\rho_{A}(x)=\gamma_{1}|\nabla u_{\theta}|^{2}+\gamma_{2}\Psi(u_{\theta}),\hskip 14.22636pt\rho_{A}(\cdot)\text{ is uniform in }t.$
where $\Psi$ is the double well potential defined by (8).
We observe that this density would be expensive to invert, thus motivating our use of the Metropolis-Hastings algorithm. With this algorithm, it is only necessary to compute the values of this density at our current and proposed sample points. Since gradients of the current sample points are already available to the network from the optimization algorithm, and the density values of the proposed sample points can be easily calculated via automatic differentiation, the algorithm can be efficiently employed in parallel to network training.
@@
Notably, we impose that the density must be uniform time. This is because if allowed to move in time, sample points will coalesce at the beginning of the time interval (as the total energy of the system decreases according to equation (5)). Exploring time dependent adaptive densities in an obvious direction to be explored in future work, and could be used to address complex time-dependent irregularities.
We then combine this with uniformly sampled points to get an effective total sampling density
(17) $\rho(x,t)=C+\alpha\rho_{A}(x)$
where $C$ is defined implicitly through $\lambda$, the proportion of points uniformly sampled as defined by (14). We denote by $\alpha$ a hyperparameter accounting for the relative importance of the adaptively sampled points to the training. This can be interpreted as weighting the relative importance for the separation of scales. With this observation, we choose this weight as exactly $\alpha=\sqrt{\gamma_{2}/\gamma_{1}}$, which dictates the separation of scales.
Points in this sampling are then weighted equally in the discrete loss computation as to truly recover the weighted loss term (13).
2.2. Metropolis-Hastings Algorithm
In order to efficiently sample from a probability distribution dependent on the network and its derivatives, we use the Metropolis-Hastings Algorithm [27, 28]. This is a sampling method which iteratively improves a collection of randomly chosen points to match the distribution given by a known probability density function. This method avoids the need to invert a distribution, and is thus ideal for settings where direct sampling is difficult.
(18) $A=\min\left(1,\frac{f(x^{\prime})g(x^{\prime}|x)}{f(x)g(x|x^{\prime})}\right)$
It is worth noting that there is a brief history of the use of Metropolis-Hastings and the broader class of Markov-Chain-Monte-Carlo (MCMC) methods in PINNs. In particular, MCMC methods are suggested by [17] as a method of implementation for residual adaptive methods when the sampling distribution is complex. Though dropout methods remain more popular in residual adaptive methods, such as was employed in [15], due to their ease of implementation. Additionally the Bayesian-PINN (B-PINN) [29] uses MCMC-based inference to capture uncertainty in the network parameters, addressing noise through probabilistic modeling rather than deterministic regularization, though the way the methods are used in this context is very different from our own.
The Metropolis-Hastings is additionally very well-studied and lends itself very naturally to the context of Physics-Informed Neural Networks. As an iteration based method, it can be run easily in conjunction with training. Additional improvements such as multiple proposal [30] can nicely address common issues in PDEs and in shocks in particular. Additionally, the method is nicely parallelizable and can be run very efficiently on GPUs [31]. These benefits make it an exceptionally good fit for sampling in the context of Neural Networks.
In our context we will use the Metropolis-Hastings method not to sample according to a likelihood or posterior distribution as is typical, but to sample according to our network-dependent heuristic function $\rho$ described above in the above Section 2 and in 2.1.
2.3. Details of Metropolis Hastings
In this section we describe in more detail the precise nature of the Metropolis-Hastings Procedure we use in our experiments. When initializing the adaptive points on each time slice, we perform $1,000$ iterations of the Metropolis-Hastings Algorithm described in 2.2. We use a proposal distribution which is a normal distribution centered around each point with standard deviation $\sqrt{\gamma_{1}}$, which is proportional to the interfacial width. We additionally update the standard deviation with each iteration to target an acceptance ratio (the proportion of proposed points accepted) of between $0.2$ and $0.6$.
The target distribution is given by the pointwise energy (17). This is easily computed by the network since the relevant gradients of existing sample points are available from the backpropagation used to compute the residual. The gradients of the proposed points can be computed in the same manner. The corrective ratio is given from the explicit formula for the pointwise density function of the normal distribution centered on each point. The acceptance probability (18) is then computed directly from these values.
We also ‘step’ the adaptive points at the conclusion of each training epoch (not after each minibatch). To do this, we perform $200$ iterations of the same algorithm described above. This ensures that the sampling distribution continuously changes to reflect changes to the solution approximation (i.e. the network).
3. Other Techniques
@@ 3.1. Time Slicing
This method was introduced in [15] and involves the gradual expansion of the time domain in discrete steps. The idea is to reduce the time complexity of the problem by first greatly restricting the time domain until it is palatable to the network. Once the problem is learned on that smaller interval, the interval can be increased. This allows for the network to only require learning a small time interval at any given time time, which is generally easier for training. This method was tested very thoroughly with a large degree of success. However, for higher-dimensional and less regular systems further issues arise.
One of the these issues is commonly refereed to as ‘Catastrophic Unlearning’ in the wider Machine Learning community, though use of this term is not commonly used in the discussion of PINNs. This issue is common in traditional data driven contexts where networks are trained to do one task and then trained to do a second separate task. In this instance the network may unlearn the original task in favor of learning the second task. In the context of PINNs this issue manifests when a network which is well trained on some sub-interval of the entire time domain, is trained on a separate or expanded time interval. As a particular example, a PINN trained on the time interval $[0,0.1]$ may drastically lose accuracy when trained on $[0,0.2]$ since simulating from 0.1 to 0.2 can be considered a ‘new task.’ Not only will the simulation not perform accurately on the new time window of 0.1 to 0.2, but it will also lose the accuracy it had on the original window of 0 to 0.1. Of course simulating accurately on $[0,0.1]$ is essential to simulating on $[0.1,0.2]$, so in the setting of PINNs the problem is even further accentuated.
The authors of [15] were aware of this issue, so they presented an alternative method that they call ‘Time Slicing II.’ This method involves the training of a separate network on each time slice rather than using the gradual expansion of training time on a single network. This does resolve the issue of catastrophic unlearning, since no network is trained to perform multiple tasks. However, it is somewhat undesirable as it requires many networks, a true time discretization, and can also dramatically increase cost in training time and number of network parameters. For these reasons we manually edit the learning rate as we evolve through time slices, which is discussed next.
@@ 3.2. Learning Rate Schedule
We additionally use a learning rate scheduler, which updates the learning rate as the training process progresses. We use a fairly simple scheduler which updates the learning rate only based off the current time slice being trained. This helps to combat the catastrophic unlearning described in the section above. We decrease the learning rate roughly linearly as the progress through the time-slices. More sophisticated methods of choosing the learning rates are constantly being studied [33, 32], however for simplicity and ease of implementation and comparison we choose only to use this simple scheme.
@@ 3.3. Residual Adaptive Sampling
A commonly employed method to increase training efficacy is Residual Adaptive sampling, first proposed in [16]. This technique involves sampling the domain in proportion to the current pointwise value of the residual. This leads to decreasing the loss very efficiently, as has been thoroughly explored in [17] and seen in many other papers including [15]. In our numerical experiments, we will use this method as a test to compare against energy adaptive sampling.
@@ 3.5. Latin Hypercube Sampling
For uniform sampling in our domain, we use Latin Hypercube sampling. This method reduces variance in sampling by partitioning the domain into a grid of $N^{d}$ hypercubes, where $N$ is the desired number of sample points and $d$ is the number of spatial dimensions. The hypercubes are then sampled randomly so that the selected hypercubes are orthogonal. This yields a selection of $N$ hypercubes in the grid, with no selected hypercube in the same row or column as any other. Once this orthogonal set of hypercubes is selected, one point within each is chosen randomly from a uniform distribution on that hypercube. These individual sample points combine to form a representative sample of $N$ points from the domain. Further explanation can be found [34].
This method guarantees that the uniform sample will be representative of the entire domain. This sampling also greatly reduces variance, and thus is less reliant on the convergence yielded by the law of large numbers. This especially helps to reduce inconsistencies of sampling too few points, and makes the training more robust.
@@ 4. Numerical Results
The following numerical benchmarks are taken from [15]. In the examples they provide, there is a clear concentration of the error in the regions of high energy. As such, this is an ideal setting for the testing of the improvements given by our new method.
For each example, we compare three methods. First, we run a high fidelity finite difference to yield an ‘exact’ solution. The second method is Residual Adaptive sampling. Here we use the Metropolis Hastings Algorithm to sample in proportion to the loss and combine these with points sampled uniformly as suggested in [17] for complex sampling distributions. Additionally this implementation uses all of the techniques described in the implementation section, including time slices, learning rate scheduling, minibatching, latin hypercube sampling, and initial condition weights. Finally we implement our energy adaptive method described in section 3.
@@
(19) $\gamma_{1}=1e-4,\hskip 5.69046pt\gamma_{2}=5,\hskip 5.69046ptu_{0}(x)=x^{2}\cos(\pi x)$
@@
Our network uses 6 fully connected layers of 128 nodes. We use the hyperbolic tangent activation function for all interior layers. For training we use $10,000$ total collocation points in the domain. We train the network until final time $t=1$, with time slices taken in increments of $0.1$. On each time slice, we run 100 epochs of ADAM training with minibatches of size $40$ split proportionally according to the adaptive point ratio, $\lambda$. As we progress through time slices, we additionally change the learning rate as described in section 3. We use learning rate $10^{-3}$ until final time 0.3, at which point we switch to $5*10^{-4}$. At final time $0.5$, we decrease to learning rate $10^{-4}$. At $0.7$ we decrease to $5*10^{-5}$, and finally at $0.9$ we decrease to $10^{-5}$.
We perform a parameter sweep on $\lambda$, the proportion of adaptively chosen points vs uniformly chosen points. The results for this are found in Figure 1. We can see from this plot that the energy adaptive method performs best with a $\lambda$ value of about $0.6$, while residual adaptive functions about equivalent for $\lambda$ between $0.6$ and $0.9$. For further discussion we consider $\lambda=0.6$ so as to more directly compare with the energy adaptive method. Notice also the consistent improvement of the energy adaptive method in both error measures.
@@
To verify that both methods are decreasing loss effectively, we display the loss across each epoch in Figure 2. It is verified here that each method is successfully evolving according to their respective losses.
@@
Figure 3. Time slices from networks which have completed the described training. We present time slices from the the residual adaptive method and the energy adaptive method at times 0, $0.25$, $0.5$, and $1$.
@@
Relative $L^{2}$ at $T=1$ $4.09\times 10^{-2}$ $1.50\times 10^{-2}$
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error $2.31\times 10^{-1}$ $7.96\times 10^{-2}$
@@
Computational Complexity
We include some brief remarks on computational cost. Our residual-based and energy-based adaptive sampling methods each required approximately $5,500$ seconds of wall-clock time, which is expected since both rely on a Metropolis–Hastings procedure with comparable proposal and acceptance steps. By comparison, a baseline PINN trained with identical hyperparameters but without adaptive sampling required approximately $3,200$ seconds. Quick computations show that the baseline method with no adaptive sampling has a cost of $3.2$ seconds per epoch, while adaptive sampling incurs an additional cost of $2.3$ seconds per epoch and $4.6$ seconds per time-slice (the cost of re-initializing the adaptive points).
While adaptive sampling therefore incurs a nontrivial additional cost, the baseline PINN completely fails to capture the solution, whereas the adaptive methods succeed. Moreover, this cost compares favorably with dropout-based residual selection strategies, which typically introduce additional overhead due to repeated residual evaluation and resampling; in practice, the total training time often increases as the fraction of retained points decreases.
@@
(20) $\gamma_{1}=1\times 10^{-4},\hskip 5.69046pt\gamma_{2}=4,\hskip 5.69046ptu_{0}(x)=x^{2}\sin(2\pi x)$
@@
Relative $L^{2}$ at $T=1$ $2.33\times 10^{-2}$ $6.87\times 10^{-3}$
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error $1.15\times 10^{-1}$ $3.20\times 10^{-2}$
@@
Finally we experiment with the 2D Example given in [15]. This is characterized by the parameters
(21) $\gamma_{1}=\lambda\epsilon^{2},\hskip 8.5359pt\gamma_{1}=\lambda,\hskip 8.5359ptu_{0}(x)=\tanh\left(\frac{.35-\sqrt{(x-.5)^{2}+(y-.5)^{2}}}{2\epsilon}\right)$
@@
Relative $L^{2}$ at $T=10$ $T=9$ $1.52\times 10^{-1}$ $3.56\times 10^{-2}$
Relative $L^{2}$ at $T=10$ $T=10$ 3.53e013.53e-01 $1.93\times 10^{-1}$
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error $T=9$ $1.73\times 10^{0}$ $9.79\times{-1}$
$L^{\infty}(0,T;L^{\infty}(\Omega))$ Error $T=10$ $1.98\times 10^{0}$ $1.90\times 10^{0}$
@@
In general there are several methods which can be used to reduce the problems posed by catastrophic unlearning. The two most prominent of which would be the employment of an xPINN [12] or the use of exponential time sampling [35]. These directions are known to be effective, but require a post-hoc knowledge of where the unlearning occurs. For our analysis here, we are trying to confine our approach only to methods which can be applied without prior knowledge of the location of the unlearning. We also believe that an intelligently chosen heuristic employed via auto-adaptive PINNs could efficiently resolve this issue. However, we leave the exploration of this as a future direction.
@@ 5. Conclusions and Future Directions
Analytic verification in the form of a proof is also desirable for the energy adaptive method, perhaps in a similar manner to the proofs in [35]. If well understood, this could also motivate choices of sampling densities for other difficult problems.